Use the given information to answer the following questions with corresponding answers.

(a)

H2G+  LaTeX: \Longleftrightarrow⟺   HG + H+ Ka1 = 4.6 LaTeX: \times× 10-3

HG  LaTeX: \Longleftrightarrow⟺   G- + H+   Ka2 = 2.5 LaTeX: \times× 10-10

Determine the dissociation constant for G- + H2O  LaTeX: \Leftrightarrow⇔   HG + OH-.

Group of answer choices

4.0 LaTeX: \times×× 10-9

4.0 LaTeX: \times×× 10-5

2.2 LaTeX: \times×× 10-12

3.5 LaTeX: \times×× 10-6

2.9 LaTeX: \times×× 10-8

(b)

Consider the following dissociation reactions of phosphate (PO43-):

PO43- + H2O  LaTeX: \Leftrightarrow⇔   HPO42- + OH-   pKb1 = 1.65

HPO42- + H2O  LaTeX: \Leftrightarrow⇔   H2PO4- + OH-   pKb2 = 6.98

H2PO4- + H2O  LaTeX: \Leftrightarrow⇔   H3PO4 + OH-   pKb3 = 11.85

Determine the dissociation constant for H2PO4-   LaTeX: \Leftrightarrow⇔   HPO42- + H+.

Group of answer choices

1.4 LaTeX: \times×× 10-12

7.1 LaTeX: \times×× 10-3

1.0 LaTeX: \times×× 10-7

4.5 LaTeX: \times×× 10-13

9.5 LaTeX: \times×× 10-8

(c)

Alanine is a diprotic acid (Ka1 = 4.57 LaTeX: \times× 10-3 and Ka2 = 2.04 LaTeX: \times× 10-10) that can have three different forms in solution: H2A+, HA, and A-. What would be the pH of the solution containing 0.085M H2A+Cl-?

Group of answer choices

3.59

2.11

4.24

2.79

3.05

(d)

Alanine is a diprotic acid (Ka1 = 4.57 LaTeX: \times× 10-3 and Ka2 = 2.04 LaTeX: \times× 10-10) that can have three different forms in solution: H2A+, HA, and A-. What would be the concentration of A- in the solution of 0.0725M H2A+Cl-?

Group of answer choices

2.04 LaTeX: \times× 10-10 M

4.57 LaTeX: \times× 10-3 M

5.38 LaTeX: \times× 10-11 M

1.48 LaTeX: \times× 10-6 M

3.15 LaTeX: \times× 10-8 M
Answer #1

(a) H2G+ <========> HG + H+             Ka1 = [HG][H+]/[H2G+] = 4.6*10-3

HG <=======> G- + H+                         Ka2 = [G-][H+]/[HG] = 2.5*10-10

Add the auto-ionization reaction of water as below.

H2O <=======> H+ + OH-                      Kw = [H+][OH-] = 1.0*10-14

The given reaction is

G- + H2O <======> HG + OH-

The equilibrium constant is written as

K = [HG][OH-]/[G-] (the concentration of H2O is assumed constant)

= ([HG]/[H+][G-])*([H+][OH-])

= Kw/Ka2

= (1.0*10-14)/(2.5*10-10)

= 4.0*10-5 (ans).

(b) The pKb values are given; determine the Kb values as

pKb = -log (Kb)

======> Kb = antilog (-pKb)

pKb

Kb

1.65

2.24*10-2

6.98

1.05*10-7

11.85

7.08*10-11

PO43- + H2O <======> HPO42- + OH-                   Kb1 = [HPO42-][OH-]/[PO43-] = 2.24*10-2

HPO42- + H2O <=======> H2PO4- + OH-              Kb2 = [H2PO4-][OH-]/[HPO42-] = 1.05*10-7

H2PO4- + H2O <=======> H3PO4 + OH-               Kb3 = [H3PO4][OH-]/[H2PO4-] = 7.08*10-11

Add the auto-ionization reaction of water as below.

H2O <=======> H+ + OH-                      Kw = [H+][OH-] = 1.0*10-14

The given reaction is

H2PO4- <=======> HPO42- + H+

The equilibrium constant is written as

K = [HPO42-][H+]/[H2PO4-]

= ([HPO42-]/[H2PO4-][OH-])*([H+][OH-])

= Kw/Kb2

= (1.0*10-14)/(1.05*10-7)

= 9.52*10-8

≈ 9.5*10-8